i need help in solving this problem
x/3-1= x/5 +1

Answer by checkley75(3666)

(X/3)-1=(X/5)+1 IF THIS IS THE PROBLEM THEN:
X/3-X/5=2
(5X-3X)/15=2 NOW CROSS MULTIPLY
5X-3X=30
2X=30
X=30/2
X=15 ANSWER.
PROOF
15/3-1=15/5+1
5-1=3+1
4=4